# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ III Improved Lower Bound on Access Cost
$\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1\leqslant n_{F}-2d-(t-1)k+2-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1)\triangleq\Delta,$
where $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1$ is defined in Lemma 5. Thus, $\Delta_{i}\leqslant\Delta$ for every $i\in[t]$. According to Lemma 6, let $\tilde{B}=\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}$ then $|\tilde{B}|=\Delta_{i}+d-1$ by $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1$. Then, we can find a $B$ satisfying condition (i)-(iii) and $|B|\geqslant(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$. This is to say that
$\Delta_{i}-|B|$ $\leqslant\Delta_{i}-(\delta-1)\left\lfloor\frac{\Delta_{i}}{r+\delta-1}\right\rfloor$
@@
$|\mathcal{R}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta\leqslant 0;\\ k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor,&\mbox{otherwise},\end{cases}$
for every $i\in[t]$. Therefore,
$\rho_{r}=\sum_{i=1}^{t}|\mathcal{R}_{i,1}|\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\\ t\left(k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),&\mbox{otherwise}.\end{cases}$
@@
Remark 3:
For the case $\delta=2$, the preceding bound is exactly the same as the one in Theorem 1 [23, Theorem IV.4]. While the case $\delta>2$, the new bound may outperform the one in Theorem 1. The reader may refer to Figure 1 as an example.
Figure 1: Compare the lower bounds on read access cost of [23, Theorem IV.4] and Theorem 2. We only consider the case $\Delta>0$ and $d\leqslant n_{I}-k+1$. Set the $t=3$ initial codes with parameters $[n_{I}=80,k=40]_{q}$, and the final code with parameters $[tk/\alpha,tk,d=15]_{q}$ and with $(r=10,\delta)_{a}$-locality where $\alpha$ is the code rate (the X-axis in the figure) of the final code. Note that the lower bound of [23, Theorem IV.4] is not relate to the $\delta$, therefore, the conditions for $\delta=3,4,5$ are the same blue circle in the figure.
@@
IV A General Construction of LRCC
In this section, we present a general construction for LRCCs with optimal access costs. The basic idea of our construction is to apply an original code to ensure a proper conversion from a set of codes, which are shortened versions of the original code. To begin, we prove the following simple lemma.
@@ IV A General Construction of LRCC
Let $g,r,\delta,\ell$ be positive integers with $g\geqslant\ell$. For an optimal $[g(r+\delta-1),\ell r,d]_{q}$-LRC $\mathcal{C}$ with $(r,\delta)_{a}$-locality, then $\mathcal{C}$ admits a parity check matrix $H$ of the form (1) where $A_{1},A_{2},\dots,A_{g}$ are $(\delta-1)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$, and $B_{1},B_{2},\dots,B_{g}$ are $((g-\ell)r)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$. Then, for any set $\mathcal{P}=\{i_{1},\dots,i_{|\mathcal{P}|}\}\subseteq[g]$ with $|\mathcal{P}|>g-\ell$, the submatrix $H||_{\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $[|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
@@
The code $\mathcal{C}$ has the parity check matrix $H$ is directly from Lemma 3. Assume that a code $\tilde{\mathcal{C}}$ has the parity check matrix $H||_{\mathcal{P}}$. On the one hand, the minimum Hamming distance $d(\tilde{\mathcal{C}})$ of $\tilde{\mathcal{C}}$ satisfies $d(\tilde{\mathcal{C}})\geqslant d$, since any $d-1$ columns of $H||_{\mathcal{P}}$ are linear independent over $\mathbb{F}_{q}$. On the other hand, the dimension $k(\tilde{\mathcal{C}})$ of $\tilde{\mathcal{C}}$ satisfies
$k(\tilde{\mathcal{C}})$ $\geqslant|\mathcal{P}|(r+\delta-1)-(|\mathcal{P}|(\delta-1)+(g-\ell)r)$
$=(|\mathcal{P}|-g+\ell)r.$
Note that $\tilde{\mathcal{C}}$ has $(r,\delta)_{a}$-locality. According to Singleton-type bound (3), we have that
$d(\tilde{\mathcal{C}})$ $\leqslant|\mathcal{P}|(r+\delta-1)-(|\mathcal{P}|-g+\ell)r+1-(|\mathcal{P}|-g+\ell-1)(\delta-1)$
@@
where the last equation holds by $\mathcal{C}$ is an optimal LRC. This is implies that $d(\tilde{\mathcal{C}})=d$, i.e., $\tilde{\mathcal{C}}$ is an optimal LRC with parameters $[|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}$. ∎
@@
Let $g,\ell,r,\delta,t$ be positive integers with $t\mid\ell$, $\ell/t>g-\ell\geqslant 0$ and $\delta\geqslant 2$. Suppose $\mathcal{C}$ is an optimal $[g(r+\delta-1),\ell r,d]_{q}$-LRC with $(r,\delta)_{a}$-locality, then there exists an access-optimal LRCC with parameters
@@
for any non-negative integer $h\leqslant g-\ell$. Moreover, the initial and final codes are optimal LRCs.
@@
Let $H$ of the form (1) be the parity check matrix of $\mathcal{C}$. Denote $s=\ell/t$. Let $\mathcal{P}_{1},\dots,\mathcal{P}_{t},\mathcal{P}$ be the partition of $[g]$ such that $|\mathcal{P}_{i}|=s$ for all $i\in[t]$ and $|\mathcal{P}|=g-\ell$. Without loss of generality, assume $\mathcal{P}_{i}=\{(i-1)s+1,\dots,is\}$ for $i\in[t]$ and $\mathcal{P}=\{\ell+1,\dots,g\}$. According to Lemma 7 and $|\mathcal{P}_{i}\cup\mathcal{P}|=s+g-\ell$, the submatrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $[(s+g-\ell)(r+\delta-1),sr,d]_{q}$-LRC $\mathcal{C}_{i}$ with $(r,\delta)_{a}$-locality for $i\in[t]$. Let $\mathcal{C}_{i}$ be the initial code with the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$. The main part of the parity check matrix concatenating is the global parity check. Therefore, we first do a puncturing operation on the initial codes.
Without loss of generality, assume $i=1$. We divide $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ into blocks as follows:
$H||_{\mathcal{P}_{1}\cup\mathcal{P}}=\left(\begin{array}[]{ccc|ccc|ccc}A_{1}&\cdots&O&O&\cdots&O&O&\cdots&O\\ \vdots&\ddots&\vdots&O&\cdots&O&O&\cdots&O\\ O&\cdots&A_{s}&O&\cdots&O&O&\cdots&O\\ \hline\cr O&\cdots&O&A_{\ell+1}&\cdots&O&O&\cdots&O\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ O&\cdots&O&O&\cdots&A_{\ell+h}&O&\cdots&O\\ \hline\cr O&\cdots&O&O&\cdots&O&A_{\ell+h+1}&\cdots&O\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ O&\cdots&O&O&\cdots&O&O&\cdots&A_{g}\\ B_{1,1}&\cdots&B_{s,1}&B_{\ell+1,1}&\cdots&B_{\ell+h,1}&B_{\ell+h+1,1}&\cdots&B_{g,1}\\ \hline\cr B_{1,2}&\cdots&B_{2,2}&B_{\ell+1,2}&\cdots&B_{\ell+h,2}&B_{\ell+h+1,2}&\cdots&B_{g,2}\end{array}\right)=\left(\begin{array}[]{c|c|c}\bm{W}_{11}&\bm{O}&\bm{O}\\ \hline\cr\bm{O}&\bm{W}_{12}&\bm{O}\\ \hline\cr\bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \hline\cr\bm{N}_{11}&\bm{N}_{12}&\bm{N}_{13}\end{array}\right),$ (15)
where $B_{i,1}$ is consisting of the first $(g-\ell-h)r$ columns of $B_{i}$, and $B_{i,2}$ is consisting of the last $hr$ columns of $B_{i}$ for $i\in[g]$.
Note that $\bm{M}_{13}$ is a $(g-\ell-h)(r+\delta-1)\times(g-\ell-h)(r+\delta-1)$ matrix. Since $\mathcal{C}_{1}$ is optimal with the minimum Hamming distance $d=\delta+(g-\ell)(r+\delta-1)$, therefore,
@@
There is an invertible square submatrix of $(\bm{M}_{13}^{\top},\bm{N}_{13}^{\top})^{\top}$ with order $(g-\ell-h)(r+\delta-1)$. Without loss of generality, let $\bm{M}_{13}$ be invertible. We derive a puncturing code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{1}$ by following elementary row operations on $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$:
$\begin{pmatrix}\bm{I}&\bm{O}&\bm{O}&\bm{O}\\ \bm{O}&\bm{I}&\bm{O}&\bm{O}\\ \bm{O}&\bm{O}&\bm{I}&\bm{O}\\ \bm{O}&\bm{O}&-\bm{N}_{13}\bm{M}_{13}^{-1}&\bm{I}\end{pmatrix}\cdot\begin{pmatrix}\bm{W}_{11}&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{12}&\bm{O}\\ \bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \bm{N}_{11}&\bm{N}_{12}&\bm{N}_{13}\end{pmatrix}=\begin{pmatrix}\bm{W}_{11}&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{12}&\bm{O}\\ \bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}&\bm{O}\end{pmatrix},$ (16)
where $\bm{I}$ is an identity matrix, $\bm{O}$ is a zero matrix.
@@
Therefore, we have that the code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ has parity check matrix
$\overline{H}_{1}=\begin{pmatrix}\bm{W}_{11}&\bm{O}\\ \bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}.$
@@
Since every $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ has the same part $H||_{\mathcal{P}}$, thus the puncturing code $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{i}$ comes from the same elementary row operations, i.e., each $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ is left-multiplied by the first matrix in the left-hand side of (16). Then, $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ has the parity check matrix $\overline{H}_{i}$ for $i\in[t]$, i.e.,
$\overline{H}_{i}=\begin{pmatrix}\bm{W}_{i1}&\bm{O}\\ \bm{O}&\bm{W}_{12}\\ \bm{N}_{i1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{i1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix},$
which holds by $\bm{j}_{12}=\bm{j}_{22}=\cdots=\bm{j}_{t2}$ for $\bm{j}\in\{\bm{W},\bm{M},\bm{N}\}$, and $\bm{z}_{13}=\bm{z}_{23}=\cdots=\bm{z}_{t3}$ for $\bm{z}\in\{\bm{M},\bm{N}\}$.

### v2

@@ III Improved Lower Bound on Access Cost
$\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1\leqslant{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{F}-2d-(t-1)k+2-\phi((t-1)k,r,\delta)}\triangleq\Delta,$
where $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$ is defined in Lemma 6. Thus, $\Delta_{i}\leqslant\Delta$ for every $i\in[t]$. According to Lemma 7, let $\tilde{V}=\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}$ then $|\tilde{V}|=\Delta_{i}+d-1$ by $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$. Then, we can find a $B$ satisfying conditions (i), (ii) and $|B|=(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$. This is to say that
$\Delta_{i}-|B|$ $=\Delta_{i}-(\delta-1)\left\lfloor\frac{\Delta_{i}}{r+\delta-1}\right\rfloor$
@@
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,&\mbox{otherwise},\end{cases}$
for every $i\in[t]$. Therefore, ρ_r = ∑_i=1^t—D_i,1—⩾{tk, if Δ⩽0,t(k-Δ+ (δ-1) ⌊Δr+δ-1⌋), otherwise.
@@
Corollary 1:
Suppose $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ corresponding to Lemma 2. Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$. Then, the write access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$ (23)
Moreover, if $d>n_{I}-k+1$, then the read access is at least $tk$.
Corollary 2:
Suppose $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$ corresponding to Lemma 3. Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $((\ell/t+h)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$ where $\delta\geqslant 2,t\mid\ell$ and $\ell/t>h\geqslant 0$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$. Then the write access cost is at least $h(r+\delta-1)$ and the read access cost is at least thrthr.
Proof.
By $\phi(k,r,\delta)=\left\lceil\frac{k}{r}\right\rceil(\delta-1)$ and Theorem 2, the write access cost $\rho_{w}$ satisfies
$\rho_{w}\geqslant n_{F}-\sum_{i\in[t]}|\mathcal{U}_{i,1}|=h(r+\delta-1)$
and
$\Delta$ $=n_{F}-2d-(t-1)k+2-\left\lceil\frac{(t-1)k}{r}\right\rceil(\delta-1)$
$=\left(\frac{\ell}{t}-h\right)(r+\delta-1)$
$>0.$
Therefore, the read access cost $\rho_{r}$ satisfies
$\rho_{r}$ $\geqslant t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right)$
$=t\left(\frac{\ell r}{t}-\left(\frac{\ell}{t}-h\right)(r+\delta-1)+\left(\frac{\ell}{t}-h\right)(\delta-1)\right)$
$=thr.$
Remark 2:
Corollary 1 extends the lower bound in Theorem 1 [27, Theorem IV.4] which is only associated with $r$, to one related to $r$ and $\delta$. For the case $\delta=2$, the new lower bound is exactly the same as the one in Theorem 1. In the case where $\delta>2$, the new lower bound on the write access cost clearly outperforms the one in Theorem 1. The expressions $\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$ are not easily reducible and LRCs achieving the minimum Hamming distance bound in Lemma 2 usually require some parameter constraints [46, 8], e.g., $r+\delta-1\mid n$, or $n=w(r+\delta-1)+m,k=ur+v$ such that $0<v<r,v+\delta-1\leqslant m<r+\delta-1$, etc. Therefore, we investigate the following specific scenarios which show the new lower bound on the read access cost surpasses the one in Theorem 1, i.e., $\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\leqslant\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$:
  1. 1.
    When $\Delta\mid(r+\delta-1)$, then
    $\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor=\frac{r\Delta}{r+\delta-1}<\frac{r\overline{\Delta}}{r+1}\leqslant\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil,$
    where the first inequality holds by $\Delta<\overline{\Delta}$ and $\delta>2$.
  2. 2.
    For $\bm{\mathcal{C}}$ with parameters $(n_{I},kr,*,*;n_{F},tkr,r,\delta)_{a}$ and the final code $\mathcal{C}_{F}$ as an optimal LRC satisfying $n_{F}=(tk+m)(r+\delta-1)$ where $m$ is a positive integer and $tk-m\geqslant 2$, then
    $\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor=(k-m)r\leqslant\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil=(k-m)r+\left\lceil\frac{r((tk-2-m)(\delta-2)-1)}{r+1}\right\rceil.$
The reader may refer to Figure 1 as an example.
Remark 3:
We would like to comment that shortly after we published our results on arXiv [14], we learned that [45] has independently obtained a similar lower bound in Corollary 1 focused on the initial codes with different parameters.
Figure 1: Compare the lower bounds on read access cost of Theorem 1  [27, Theorem IV.4] and Corollary 1. We consider only the case $\Delta>0$ and $d\leqslant n_{I}-k+1$. Set the $t=3$ initial codes with parameters $n_{I}=80,k=40]_{q}$, and the final code with parameters $tk/\alpha,tk,d=15]_{q}$ and with $(r=10,\delta)_{a}$-locality where $\alpha$ is the code rate (the X-axis in the figure) of the final code. Note that the lower bound of Theorem 1 is independent of the $\delta$, therefore, the conditions for $\delta=3,4,5$ are the same blue circle in the figure.
IV A General Construction of LRCCs
In this section, we present a general construction of LRCCs with optimal access costs and optimal initial and final codes which achieve the minimum Hamming distance bound in (3). The basic idea of our construction is to choose the suitable initial and final codes from the punctured codes of a given base code. To begin, we prove the following simple lemma.
@@ IV A General Construction of LRCCs
Let $g,r,\delta,\ell$ be positive integers with $g\geqslant\ell$.
  1. 1.
    For an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC $\mathcal{C}$ with $(r,\delta)_{a}$-locality and achieving the minimum Hamming distance bound in (3), the $\mathcal{C}$ admits a parity check matrix $H$ of the form (1) where $A_{1},A_{2},\dots,A_{g}$ are $(\delta-1)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$, and $B_{1},B_{2},\dots,B_{g}$ are $((g-\ell)r)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$.
  2. 2.
    For any set $\mathcal{P}=\{i_{1},\dots,i_{|\mathcal{P}|}\}\subseteq[g]$ with $|\mathcal{P}|>g$-$\ell$, the submatrix $H||_{\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}-$LRC with $(r,\delta)_{a}$-locality and achieving the minimum Hamming distance bound in (3).
@@
[…]
$\geqslant|\mathcal{P}|(r+\delta-1)-(|\mathcal{P}|(\delta-1)+(g-\ell)r)$
$=(|\mathcal{P}|-g+\ell)r,$
where the first inequality holds by the matrix $H||_{\mathcal{P}}$ has $|\mathcal{P}|(\delta-1)+(g-\ell)r$ rows. Note that $\tilde{\mathcal{C}}$ has $(r,\delta)_{a}$-locality. According to Singleton-type bound (3), we have that
[…]
@@
where the last equation holds by $\mathcal{C}$ is an optimal LRC. This implies that $d(\tilde{\mathcal{C}})=d$, i.e., $\tilde{\mathcal{C}}$ is an optimal LRC with parameters $|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}$. ∎
@@
Let $g,\ell,r,\delta,t$ be positive integers such that $t\mid\ell$, $\ell/t>g-\ell\geqslant 0$ and $\delta\geqslant 2$. Suppose $\mathcal{C}$ is an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC with $(r,\delta)_{a}$-locality. Then there exists an access-optimal LRCC with parameters
@@
for any non-negative integer $h\leqslant g$-$\ell$. Moreover, the initial and final codes achieve the minimum Hamming distance bound in (3).
@@
Choose the initial codes: Let $H$ be the parity check matrix of $\mathcal{C}$, where $H$ is of the form given in (1). Denote $s=\ell/t$. Let $\mathcal{P}_{1},\dots,\mathcal{P}_{t},\mathcal{P}$ be a partition of $[g]$ such that $|\mathcal{P}_{i}|=s$ for all $i\in[t]$ and $|\mathcal{P}|=g$-$\ell$. Without loss of generality, assume $\mathcal{P}_{i}=\{(i-1)s+1,\dots,is\}$ for $i\in[t]$ and $\mathcal{P}=\{\ell+1,\dots,g\}$. According to Lemma 8 and $|\mathcal{P}_{i}\cup\mathcal{P}|=s+g$-$\ell$, the submatrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $(s+g-\ell)(r+\delta-1),sr,d]_{q}-$LRC $\mathcal{C}_{i}$ with $(r,\delta)_{a}$-locality for $i\in[t]$. Let $\mathcal{C}_{i}$ be the initial code with the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$.
Puncture the initial codes: Without loss of generality, assume $i=1$. We divide $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ into blocks as follows:
$H||_{\mathcal{P}_{1}\cup\mathcal{P}}=\left(\begin{array}[]{ccc|ccc|ccc}A_{1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&\cdots&A_{s}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&A_{\ell+1}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{\ell+h}&\bm{0}&\cdots&\bm{0}\\ \hline\cr\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&A_{\ell+h+1}&\cdots&\bm{0}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&\bm{0}&\bm{0}&\cdots&A_{g}\\ B_{1,1}&\cdots&B_{s,1}&B_{\ell+1,1}&\cdots&B_{\ell+h,1}&B_{\ell+h+1,1}&\cdots&B_{g,1}\\ \hline\cr B_{1,2}&\cdots&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{s,2}}&B_{\ell+1,2}&\cdots&B_{\ell+h,2}&B_{\ell+h+1,2}&\cdots&B_{g,2}\end{array}\right)\triangleq\left(\begin{array}[]{c|c|c}W_{11}&\bm{0}&\bm{0}\\ \hline\cr\bm{0}&W_{12}&\bm{0}\\ \hline\cr M_{11}&M_{12}&M_{13}\\ \hline\cr N_{11}&N_{12}&N_{13}\end{array}\right),$ (25)
where $B_{i,1}$ consists of the first $(g-\ell-h)r$ rows of $B_{i}$, and $B_{i,2}$ consists of the last $hr$ rows of $B_{i}$ for $i\in[g]$.
Since $\mathcal{C}_{1}$ is optimal with the minimum Hamming distance $d=\delta+(g-\ell)(r+\delta-1)$, therefore,
@@
There is an invertible square submatrix of $(M_{13}^{\top},N_{13}^{\top})^{\top}$ with order $(g-\ell-h)(r+\delta-1)$. Note that $M_{13}$ is a $(g-\ell-h)(r+\delta-1)\times(g-\ell-h)(r+\delta-1)$ matrix, and rank( (Aℓ+h+10 ⋮⋱⋮0 ⋯Ag) ) = (g-ℓ-h)(δ-1).
By exchanging the last $(g-\ell)r$ rows of $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$, the matrix $M_{13}$ can be invertible. We derive a punctured code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{1}$ by performing elementary row operations on $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$:
$\begin{pmatrix}I&\bm{0}&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&-N_{13}M_{13}^{-1}&I\end{pmatrix}\cdot\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}&N_{12}&N_{13}\end{pmatrix}=\begin{pmatrix}W_{11}&\bm{0}&\bm{0}\\ \bm{0}&W_{12}&\bm{0}\\ M_{11}&M_{12}&M_{13}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&N_{12}-N_{13}M_{13}^{-1}M_{12}&\bm{0}\end{pmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\triangleq H_{1}^{*}}.$ (26)
In fact that the dual of puncturing code $\mathcal{C}_{1}$ is equivalent to the shortening of the dual code $\mathcal{C}_{1}^{\bot}$, and
$(\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]})^{\bot}=\left\{\bm{c}|_{[(s+h)(r+\delta-1)]}\middle|\begin{array}[]{c}\bm{c}\in\bm{m}H_{1}^{*},\bm{m}\in\mathbb{F}_{q}^{s(\delta-1)+(g-\ell)(r+\delta-1)},\\ \bm{c}|_{[(s+g-\ell)(r+\delta-1)]\setminus[(s+h)(r+\delta-1)]}=\bm{0}\end{array}\right\}.$
@@
Therefore, the punctured code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ from (26) has the parity check matrix
$\overline{H}_{1}=\begin{pmatrix}W_{11}&\bm{0}\\ \bm{0}&W_{12}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}$.
@@
Since every $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ has the same part $H||_{\mathcal{P}}$, thus the punctured code $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{i}$ comes from the same elementary row operations, i.e., each $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ is left-multiplied by the first matrix in the left-hand side of (26). Then, $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ has the parity check matrix $\overline{H}_{i}$ for $i\in[t]$, i.e.,
$\overline{H}_{i}=\begin{pmatrix}W_{i1}&\bm{0}\\ \bm{0}&W_{12}\\ N_{i1}-N_{13}M_{13}^{-1}M_{i1}&N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}$,
which holds by $j_{12}=j_{22}=\cdots=j_{t2}$ for $j\in\{W,M,N\}$, and $z_{13}=z_{23}=\cdots=z_{t3}$ for $z\in\{M,N\}$.